# Avatar Concept Slider: Manipulate Concepts In Your Human Avatar With Fine-grained Control

### v2

@@ Abstract
Language based editing of 3D human avatars to precisely match user requirements is challenging due to the inherent ambiguity and limited expressiveness of natural language. To overcome this, we propose the Avatar Concept Slider (ACS), a 3D avatar editing method that allows precise manipulation of semantic concepts in human avatars towards a specified intermediate point between two extremes of concepts, akin to moving a knob along a slider track. To achieve this, our ACS has three designs. 1) A Concept Sliding Loss based on Linear Discriminant Analysis to pinpoint the concept-specific axis for precise editing. 2) An Attribute Preserving Loss based on Principal Component Analysis for improved preservation of avatar identity during editing. 3) A 3D Gaussian Splatting primitive selection mechanism based on concept-sensitivity, which updates only the primitives that are the most sensitive to our target concept, to improve efficiency. Results demonstrate that our ACS enables fine-grained 3D avatar editing with efficient feedback, without harming the avatar quality or compromising the avatar’s identifying attributes.
@@ 1 Introduction
Creating high-fidelity human avatars and editing them according to user demands has shown its importance in multiple scenarios such as game development, film production and virtual character creation for the metaverse and live streaming applications [1, 42, 39]. To meet this demand, many works [16, 20, 19, 24, 47, 26, 4] have made significant progress in generating 3D human avatars with simple yet user-friendly language prompts, demonstrating a strong ability to generate high-fidelity and realistic avatars. At the same time, there has also been an increased demand for better control over the creation and customization of personalized digital avatars [46], e.g., changing the hair style, clothing, or adding some accessories to the avatar. To achieve this while avoiding the time-consuming process of repeated generation and prompt modifications, avatar editing techniques, which aim to edit and sculpt some target details of an avatar while preserving most of the original identifying information, have emerged as a promising paradigm.
Due to its importance, avatar editing has attracted much research attention recently [46, 30, 2, 11]. Drawing inspiration from the successes achieved in harnessing diffusion models for 3D content generation [34, 25, 40], previous works such as HeadSculpt [11] edit avatars by leveraging an instruction-guided diffusion model [3]. Likewise, TECA [46] performs localized editing using a text-driven diffusion model with guidance from a 2D segmentation model [27].
Despite the significant progress, existing 3D avatar editing methods [46, 11] are limited due to their reliance on text prompts as the sole guidance signals. Specifically, text prompts can be quite ambiguous, making it difficult to edit the avatar to precisely align with the users’ expectations. For instance, to describe the desired hair length of a human avatar, users may prefer descriptive words such as ‘long’ or ‘short’ – which are ambiguous and limited in expressing the exact degree of length. Moreover, this difficulty in achieving precise editing with text alone becomes even more evident when manipulating complex concepts like ‘sharpness’ of facial features, or other adjectives which may be hard to put into words, e.g., ‘kind’ or ‘evil’. In these cases, multiple aspects of the avatar need to be simultaneously considered, such as the shape of each facial feature, body posture, hairstyle and so on, which are very challenging to be clearly expressed in one or even a few text prompts, much less allowing precise control over them.
@@
Figure 1: Illustration of our fine-grained concept editing. In this example, the user has provided text descriptions for two opposing ends of a concept: ‘slender’ vs ‘muscular’. Our ACS allows users to precisely specify the exact level of concept expression (e.g., amount of muscle) that is desired, by moving the knob on the slider bar.
@@
Driven by the problems mentioned above, our aim in this work is to precisely edit target concepts of avatars in a convenient manner, akin to moving the knob on a continuous slider bar, from which users may select exact values along the bar to specify their desired avatar. An overview is shown in Fig. 1. However, achieving this aim is challenging, mainly due to three reasons: 1 For each concept (e.g., body shape), it can be challenging to find such a continuous bar or axis to express the intermediate degree of avatar concepts, since these concepts are often high-level abstract concepts that exist beyond the pixel level. This makes it difficult to identify where these high-level concepts reside in feature space, as well as pinpoint the precise axis that links the two opposing ends of a specific concept within the large feature space. 2 Furthermore, it is challenging to isolate and edit only the desired concept without changing other identifying information of the avatar. This is because each avatar is generated with a mixture of elements, such as face shape, hairstyle, and clothing type, which are also entangled with concepts such as gender, race, and age. While editing the target concept, we may inevitably alter other attributes that we want to retain, and it is challenging to prevent this from happening. 3 Lastly, while achieving the slider-like fine-grained control, the editing process should also ideally be quick, to provide quick feedback so that users can iteratively refine the avatar towards the desired degree of concept expression (e.g., age of the avatar) by adjusting the slider knob. This puts high demands on the rendering and editing speed.
@@
In this work, we propose a novel Avatar Concept Slider (ACS) for 3D human avatar editing, which enables users to edit their 3D avatar precisely towards the desired degree of expression of a given concept, offering much more fine-grained manipulation of concepts than with text inputs alone. To overcome the three challenges highlighted above, our ACS comprises three corresponding designs, as follows: 1 To pinpoint the precise concept axis that links the two opposing ends of a specific concept, we introduce the Concept Sliding Loss based on Linear Discriminant Analysis (LDA) to fine-tune an adapter. This facilitates precise slider-like editing using the score distillation sampling (SDS) optimization editing pipeline. LDA facilitates the identification of a continuous 1D axis (which we call the concept axis) that is the most discriminative in linking the two opposing sides of the user-provided concept. 2 To preserve key identifying attributes during editing, we further devise an Attribute Preserving Loss based on Principal Component Analysis (PCA), leveraging PCA to extract key attribute information from the features that are orthogonal to the target concept. Then, by encouraging these key attribute information to be retained, it enables users to edit their target concept in isolation, disentangled from other identifying attributes. 3 For higher efficiency in rendering and optimization, we employ the 3D Gaussian Splatting (3DGS) representation for our avatar during editing. Moreover, we further propose a 3DGS primitive selection mechanism based on their concept-sensitivity, which enables us to selectively optimize a small set of the most crucial Gaussian primitives to reduce redundancy and improve efficiency. With the above designs incorporated into the deployed ACS, users can input descriptions of two opposing ends of a target concept, e.g., with a pair of text phrases. Subsequently, users can finely adjust the degree of expression of a desired concept on a given avatar, by simply moving a knob across a slider bar.
@@ 2 Related Work
3D Gaussian Splatting (3DGS). Recently introduced by [23], 3DGS is a 3D representation that has attracted much attention across various domains due to its fast rendering speed. For instance, some works explore using 3DGS to model dynamic scenes [41, 28] or for SLAM [18, 22], leading to improved rendering speed and visual quality in both cases. Moreover, some works also explore using 3DGS for generative models, e.g., text-to-3D content generation [37, 45, 6], or human avatar generation [51, 26]. In this work, we employ 3DGS due to its efficiency and quality. To further improve the rendering speed of 3DGS for our avatar editing task, we introduce a mechanism to selectively edit only the primitives that are the most relevant to the target concept.
Text-Driven 3D Human Avatar Generation. To streamline the labour-intensive 3D avatar creation process, many recent works [16, 20, 48, 47, 4, 24, 19] focus on text-driven 3D avatar generation, relying solely on text prompt inputs to create a desired high-fidelity 3D human avatar. One line of works optimize their avatar model with text guidance in CLIP [35] space, e.g., AvatarCLIP [16]. More recent works such as DreamHuman [24], DreamAvatar [4], DreamWaltz [19], and AvatarVerse [47] rely on a text-to-image 2D diffusion prior (e.g., [36, 49]) to optimize their 3D avatar via an SDS optimization pipeline [34]. More recently, HumanGaussian [26] proposes to use 3DGS to represent the human avatar, resulting in faster convergence and high-quality avatars. However, these text-driven avatar generation methods are limited by the expressivity and ambiguity that is inherent in language, and often struggle with precisely generating the user-desired avatars with text prompts alone.
Text-Driven 3D Human Avatar Editing. Besides generation, some works [30, 46, 2] focus on editing existing 3D human avatars, sculpting the avatars according to the desired shape and texture. AvatarStudio [30] presents a text-driven method to edit head avatars based on the neural volumetric scene representation [38]. TECA [46] leverages a segmentation model to achieve localized editing based on text prompts. Another related line of works [12, 50, 5] further investigates the editing of entire 3D scenes via text instructions. Despite impressive generation and editing results, these methods are limited by their reliance on language instructions, which can be ambiguous regarding the degree of expression of the target concept. Beyond these works, this paper aims at improving avatar editing methods with finer-grained control, enabling users to manipulate the avatar more precisely with respect to a target concept.
@@ 3 Background: SDS Optimization Pipeline
Score Distillation Sampling (SDS) [34] leverages 2D diffusion models as a strong prior for 3D representation learning, e.g., text-to-3D content generation [25, 20] and 3D avatar editing [30]. Given a 3D representation $\boldsymbol{\theta}$ (e.g., NeRF [31], 3DGS [23]), SDS optimization pushes the renderings of the 3D representation towards the target distribution (e.g., high-quality natural images), by distilling the predicted score from a pre-trained 2D diffusion prior. By using text-to-image diffusion models [36] as a prior, we can facilitate text-to-3D generation by optimizing the 3D representation via SDS.
@@
Overall, this mechanism allows us to efficiently adapt a diffusion model with a small set of parameters $\Delta\boldsymbol{\phi}$ to cater for specific requirements, e.g., for editing, or for using additional control information. Hence, we also adopt the SDS optimization and LoRA adapter pipeline as introduced in Eq. 1.
@@ 4 Method: Avatar Concept Slider
Figure 2: (a) Overview of the fine-tuning stage, where an adapter is fine-tuned to learn slider-like capabilities. Firstly, using the provided descriptions of the positive ($c_{p}$) and negative ($c_{n}$) side of the concept, we extract the corresponding positive and negative features. Then, to fine-tune the adapter $\Delta\phi$, our proposed Concept Sliding Loss (Sec. 4.1) is applied to learn the ability to slide across opposing ends of the target concept. The attribute-preserving loss (Sec. 4.2) helps to retain the key identifying attributes of the avatar. (b) After fine-tuning, the adapter is applied in an SDS optimization pipeline to achieve fine-grained concept-specific 3D avatar editing. For fast rendering and optimization speeds, we leverage 3DGS representation for our 3D avatar. The proposed primitive selection mechanism (Sec. 4.3) further improves efficiency, by optimizing only the most related primitives to the target concept.
To customize a human avatar, previous text-driven 3D avatar editing methods [30, 46, 11] are limited in terms of precision. To overcome this limitation, this paper aims to achieve precise and fine-grained manipulation of 3D avatars towards a target concept in a convenient manner, akin to sliding a knob along a slider bar. However, achieving this goal is challenging due to the three reasons outlined in the introduction. To address the three challenges, we propose our Avatar Concept Slider (ACS), discussed in the following subsections. In Sec. 4.1, we address challenge 1 by leveraging LDA and devising a sliding loss to enable slider-like fine-grained control of the editing process. To tackle challenge 2, discussed in Sec. 4.2, we augment the sliding loss with an attribute-preserving loss based on PCA, ensuring concept-specific editing that is disentangled from the identifying attributes to be preserved. Lastly, in Sec. 4.3, we introduce a concept-sensitivity-based primitive selection mechanism, facilitating faster editing via selective optimization of our avatar’s 3DGS primitives, for handling challenge 3. Overall, our pipeline is built upon SDS optimization (introduced in Sec. 3), and an illustration of our full method is shown in Fig. 2.
@@ 4.1 Precise Slider-like Concept Editing
To precisely control the 3D avatar editing towards a specific degree of the target concept, the first challenge is to identify a precise and continuous axis connecting two opposing extremes of the concept. In particular, this can be challenging because the concepts are often abstract and high-level, making it difficult to pinpoint where exactly these opposing sides and their connecting axis reside in the large feature space. To overcome this, we leverage LDA to identify a continuous concept axis between the user-provided pair of descriptions for the two opposing sides, directly computing a 1D axis that is the most discriminative to link the opposing sides within the large feature space. Then, based on the identified concept axis, we introduce a Concept Sliding Loss ($\mathcal{L}_{\text{sliding}}$) to fine-tune an adapter, which enables fine-grained concept-specific editing via SDS optimization (Sec. 3). Below, we first describe how we identify a concept axis, before discussing the Concept Sliding Loss.
@@
Concept Axis Identification with LDA. First, we aim to identify a continuous concept axis between the user-provided pair of descriptions that represent two opposing sides of the target concept, where each point on the axis represents an intermediate degree of expression of the concept. By identifying the concept axis, we can precisely control the transitioning of the concept from one end to the other, and facilitate the learning of sliding capability via the sliding loss, enabling users to select any intermediate point along the concept axis for precise editing. Yet, this task is not straight forward. For instance, a simple solution may be to interpolate the semantic embeddings of two concept extremes, but such simple interpolation operations in a high-dimensional space often incurs high information redundancy and concept entanglement, incorporating many other unrelated concepts.
@@
Therefore, to handle this challenge, we leverage two-class LDA. Specifically, LDA [13, 29] is a statistical technique that aims to identify the crucial information that can well-separate different groups/classes of data samples, even while using a low dimensional representation. Thus, by employing two-class LDA, we are able to directly compute a 1D axis (resembling a two-sided slider bar) that is the most crucial and discriminative to link the opposing concepts (and their corresponding sets of features) within the large feature space. More precisely, we harness LDA to exploit the latent space of the pre-trained text-to-image diffusion model used during SDS optimization (Sec. 3) to facilitate understanding of the target concept in a high-quality and consistent latent space.
@@
Here, we present the details of how we identify a concept axis with two-cla ss LDA. At the start, the user provides a pair of text prompts representing the positive ($c_{p}$) and negative ($c_{n}$) sides of a concept. Then, based on these prompts, we sample two sets of latent features from the diffusion model $\boldsymbol{\epsilon}_{\boldsymbol{\phi}}$ by running the generation process $N_{s}$ times for each of the positive and negative sides, where we perform sampling over different diffusion timesteps using different seeds. This yields two sets of features $\{\mathbf{F}_{p,i,x,y}\}_{i\in[1,N_{s}],x\in[1,H],y\in[1,W]}$ and $\{\mathbf{F}_{n,i,x,y}\}_{i\in[1,N_{s}],x\in[1,H],y\in[1,W]}$ corresponding to the positive and negative sides, where $H,W$ refer to the spatial dimensions of latent features. Each feature vector is $D$-dimensional, i.e., $\mathbf{F}_{p,i,x,y}\in\mathbb{R}^{D}$, $\mathbf{F}_{n,i,x,y}\in\mathbb{R}^{D}$. For clarity, we do not incorporate the diffusion time step in this formulation.
@@
We then start the LDA analysis [13, 29] using these two feature sets for the two opposing sides of the target concept. Specifically, we first compute both the within-class scatter ($\mathbf{S}_{w}\in\mathbb{R}^{D\times D}$) and the between-class scatter ($\mathbf{S}_{b}\in\mathbb{R}^{D\times D}$), where $\mathbf{S}_{w}$ quantifies how much individual feature vectors deviate from their corresponding set centroids, while $\mathbf{S}_{b}$ measures the deviation between the centroids of both sides of the concept (more details in Supplementary). Then, using $\mathbf{S}_{b}$ and $\mathbf{S}_{w}$, the goal of LDA is to find a concept axis $\mathbf{b}_{c}\in\mathbb{R}^{D}$, a 1D axis such that the projection of each feature vector onto the concept axis $\mathbf{b}_{c}$ leads to maximal ratio of $\mathbf{S}_{b}$ to $\mathbf{S}_{w}$. Intuitively, this encourages the feature vectors from both sets to be separated and positioned far from each other along the concept axis, while the feature vectors from the same set become tightly clustered along the concept axis. By doing so, LDA effectively pinpoints the direction of the 1D concept axis, along which the opposing ends of the concept reside. Specifically, this process can be expressed as the following optimization problem:
@@
where the numerator ($\mathbf{w}\mathbf{S}_{b}\mathbf{w}^{T}$) and denominator ($\mathbf{w}\mathbf{S}_{w}\mathbf{w}^{T}$) on the right hand side of Eq. 2 represent the projected scatters over the concept axis, and we want to find a concept axis $\mathbf{b}_{c}$ that maximises their ratio while constraining $\mathbf{w}$ to be a unit vector. Notably, the solution to Eq. 2 can be conveniently computed as the leading eigenvector of $\mathbf{S}_{w}^{-1}\mathbf{S}_{b}$ (refer to Supplementary).
@@
Concept Sliding Loss. The above process identifies a concept axis which represents the continuous transition between the two opposing sides of the concept under consideration. Based on the identified concept axis, we introduce a Concept Sliding Loss to fine-tune an adapter to facilitate concept-specific editing of the 3D avatar. Specifically, since we adopt the SDS optimization pipeline (Sec. 3) to edit the 3D avatar via a pre-trained diffusion model, we aim to incorporate such concept-specific slider-like editing abilities into it. Hence, we propose a Concept Sliding Loss ($\mathcal{L}_{\text{sliding}}$) to impart such concept transitioning abilities to the adapter, enabling us to precisely control the editing through a scale factor $\alpha$ in Eq. 1.
@@
Intuitively, our Concept Sliding Loss $\mathcal{L}_{sliding}$ facilitates the controlled movements of features along the concept axis towards a specified intermediate point, which is specified by the user as a slider interpolation factor. Specifically, to control such movements of features during editing, our loss fine-tunes the adapter ($\Delta\phi$) such that the scale factor $\alpha$ in Eq. 1 gains the ability as the slider interpolation factor. Thus, at deployment, with the user-provided interpolation factor $\hat{\alpha}$, we can set it as the scale factor (i.e., setting $\alpha=\hat{\alpha}$), thus allowing the adapted diffusion model to precisely locate the desired intermediate point on the concept axis. The Concept Sliding Loss $\mathcal{L}_{sliding}$ is defined as:

### v3

# Avatar Concept Slider: Controllable Editing of Concepts in 3D Human Avatars

@@ Abstract
Text-based editing of 3D human avatars to precisely match user requirements is challenging due to the inherent ambiguity and limited expressiveness of natural language. To overcome this, we propose the Avatar Concept Slider (ACS), a 3D avatar editing method that allows precise editing of semantic concepts in human avatars towards a specified intermediate point between two extremes of concepts, akin to moving a knob along a slider track. To achieve this, our ACS has three designs: Firstly, a Concept Sliding Loss based on linear discriminant analysis to pinpoint the concept-specific axes for precise editing. Secondly, an Attribute Preserving Loss based on principal component analysis for improved preservation of avatar identity during editing. We further propose a 3D Gaussian Splatting primitive selection mechanism based on concept-sensitivity, which updates only the primitives that are the most sensitive to our target concept, to improve efficiency. Results demonstrate that our ACS enables controllable 3D avatar editing, without compromising the avatar quality or its identifying attributes.
@@ 1 Introduction
Creating high-fidelity human avatars and editing them according to user demands has shown its importance in multiple scenarios such as game development, film production and virtual character creation for the metaverse and live streaming applications [1, 37, 35]. To meet this demand, many works [12, 15, 14, 19, 39, 21, 4] have made significant progress in generating 3D human avatars with simple yet user-friendly language prompts, demonstrating a strong ability to generate high-fidelity and realistic avatars. At the same time, there has also been an increased demand for better control over the creation and customization of personalized digital avatars [40], e.g., changing the hair style, clothing, or adding some accessories to the avatar. To achieve this while avoiding the time-consuming process of repeated generation and prompt modifications, avatar editing techniques have emerged as a promising paradigm. These techniques aim to edit and sculpt some target details of an avatar while preserving most of the original identifying information.
@@
Figure 1: Illustration of our controllable concept editing. In this example, the user has provided text descriptions for two opposing ends of a concept: ‘short sleeves’ vs ‘long sleeves’. Our ACS allows users to specify the exact level of concept expression (e.g., sleeve length) that is desired, by moving the knob on the slider bar.
Due to its importance, avatar editing has attracted much research attention recently [40, 24, 2, 9]. Drawing inspiration from the successes in harnessing diffusion models for 3D content generation [28, 20, 36], previous works such as HeadSculpt [9] edit avatars by using an instruction-guided diffusion model [3]. Likewise, TECA [40] performs localized editing using a text-driven diffusion model with guidance from a segmentation model [22]. Overall, these methods have led to advancements in text-based 3D avatar editing.
Despite the significant progress, existing 3D avatar editing methods [40, 9] are limited due to their reliance on text prompts as the sole guidance signals. Specifically, text prompts can be quite ambiguous, making it difficult to edit the avatar to precisely align with the users’ expectations. For instance, to describe the desired hair length of a human avatar, users may prefer descriptive words such as ‘long’ or ‘short’ – which are ambiguous and limited in expressing the exact degree of length. Moreover, this difficulty in achieving precise editing with text alone becomes even more evident when manipulating complex concepts like ‘sharpness’ of facial features, or other adjectives which may be hard to put into words, e.g., ‘kind’ or ‘evil’.
@@
Driven by the problems mentioned above, our aim is to precisely edit target concepts of avatars in a convenient manner, akin to moving the knob on a continuous slider bar, from which users may select exact values along the bar to specify their desired avatar. An overview is shown in Fig. 1. However, achieving this aim is challenging, because: 1 For each concept (e.g., body shape), it can be challenging to find such a continuous bar to express the intermediate degree of avatar concepts, since these concepts are often high-level abstract concepts that exist beyond the pixel level. This makes it difficult to identify where these high-level concepts reside in feature space, as well as pinpoint the precise axes that link the two opposing ends of a specific concept within the large feature space. 2 Furthermore, it is challenging to isolate and edit only the desired concept without changing other identifying information of the avatar. This is because each avatar is generated with a mixture of elements, such as face shape, hairstyle, and clothing type, which are also entangled with concepts such as gender, race, and age. While editing the target concept, we may inevitably alter other attributes that we want to retain, and it is challenging to prevent this from happening.
@@
In this work, we propose a novel Avatar Concept Slider (ACS) for 3D human avatar editing, which enables users to edit their 3D avatar precisely towards the desired degree of expression of a given concept, offering much more controllable editing of concepts than with text inputs alone. To overcome the three challenges highlighted above, our ACS comprises three corresponding designs, as follows: 1 To pinpoint the precise concept axes that link the two opposing ends of a specific concept, we introduce the Concept Sliding Loss based on Linear Discriminant Analysis (LDA) to fine-tune an adapter. This facilitates precise slider-like editing using the score distillation sampling (SDS) optimization editing pipeline. LDA facilitates the identification of continuous axes (which we call the concept axes) that are the most discriminative in linking the two opposing sides of the user-provided concept. 2 To preserve key identifying attributes during editing, we further devise an Attribute Preserving Loss based on Principal Component Analysis (PCA), leveraging PCA to extract key attribute information from the features that are orthogonal to the target concept. Then, by encouraging these key attribute information to be retained, it enables users to edit their target concept in isolation, disentangled from other identifying attributes. 3 Moreover, we further propose a 3DGS primitive selection mechanism based on their concept-sensitivity, which enables us to selectively optimize a small set of the most crucial Gaussian primitives to reduce redundancy and improve efficiency.
With the above designs incorporated into the deployed ACS, users can input descriptions of two opposing ends of a target concept, e.g., with a pair of text phrases. Subsequently, users can precisely adjust the degree of expression of a desired concept on a given avatar, by simply moving a knob across a slider bar, achieving controllable editing.
@@ 2 Related Work
Text-Driven 3D Human Avatar Generation. To streamline the labour-intensive 3D avatar creation process, many recent works [12, 15, 41, 39, 4, 19, 14] focus on text-driven 3D avatar generation, relying solely on text prompt inputs to create a desired high-fidelity 3D human avatar. One line of works optimize their avatar model with text guidance in CLIP [29] space, e.g., AvatarCLIP [12]. Subsequent works such as DreamHuman [19], DreamAvatar [4], DreamWaltz [14], AvatarVerse [39], and TEDRA [32] rely on a text-to-image 2D diffusion prior (e.g., [30, 42]) to optimize their 3D avatar via an SDS optimization pipeline [28]. Recently, HumanGaussian [21] proposes to use 3DGS to represent the human avatar, resulting in faster convergence and high-quality avatars. However, these text-driven avatar generation methods are limited by the expressivity and ambiguity that is inherent in language, and often struggle with precisely generating the user-desired avatars with text prompts alone.
Text-Driven 3D Human Avatar Editing. Besides generation, some works [24, 40, 2] focus on editing existing 3D human avatars, sculpting the avatars according to the desired shape and texture. AvatarStudio [24] presents a text-driven method to edit head avatars based on the neural volumetric scene representation [34]. TECA [40] leverages a segmentation model to achieve localized editing based on text prompts. Another related line of works [10, 45, 5] further investigates the editing of entire 3D scenes via text instructions. Despite impressive generation and editing results, these methods are limited by their reliance on language instructions, which can be ambiguous regarding the degree of expression of the target concept. Orthogonal to these works, our paper aims at improving the controllability of 3D avatar editing methods, enabling users to edit the avatar more precisely towards the desired degree of a target concept.
Controllable Editing. Controllable editing methods are increasingly crucial for enhancing generative flexibility in real-world applications. In the image domain, ControlNet [43] is a representative work which enables the conditioning of image generation on different controls, beyond text-based inputs. Subsequent works [16, 44, 7] explore this further, controlling the spatial layouts [44], injecting different conditions [16], or editing specific concepts in images [7]. Differently, we are the first to explore controllable 3D human avatar editing, by precisely controlling their concepts via a slider. Our proposed ACS enables precise control via a slider, which is integrated into the SDS distillation pipeline and adjusts the concepts distilled into the 3D avatar. To achieve controllability and identity preservation, we also introduce an LDA-based Concept Sliding Loss and PCA-based Attribute Preserving Loss.
@@ 3 Background: SDS Optimization Pipeline
Score Distillation Sampling (SDS) [28] leverages 2D diffusion models as a strong prior for 3D representation learning, e.g., text-to-3D content generation [20, 15] and 3D avatar editing [24]. Given a 3D representation $\boldsymbol{\theta}$ (e.g., NeRF [25], 3DGS [18]), SDS optimization pushes the renderings of the 3D representation towards the target distribution (e.g., high-quality natural images), by distilling the predicted score from a pre-trained 2D diffusion prior. Thus, by using text-to-image diffusion models [30], we can facilitate text-to-3D generation by optimizing the 3D representation via SDS.
@@
Overall, this mechanism allows us to efficiently adapt a diffusion model with a small set of parameters $\Delta\boldsymbol{\phi}$ to cater for specific requirements, e.g., for editing, or for using additional control information. Hence, we also adopt the SDS optimization and LoRA adapter pipeline in Eq. 1.
@@ 4 Method: Avatar Concept Slider
Figure 2: (a) Overview of the fine-tuning stage, where an adapter is fine-tuned to learn slider-like capabilities. Firstly, using the provided descriptions of the positive ($c_{p}$) and negative ($c_{n}$) side of the concept, we extract the corresponding positive and negative features. Then, to fine-tune the adapter $\Delta\phi$, our proposed Concept Sliding Loss (Sec. 4.1) is applied to learn the ability to slide across opposing ends of the target concept. The attribute-preserving loss (Sec. 4.2) helps to retain the key identifying attributes of the avatar. (b) After fine-tuning, the adapter is applied in an SDS optimization pipeline to achieve controllable concept-specific 3D avatar editing. The proposed 3DGS primitive selection mechanism (Sec. 4.3) further improves efficiency, by optimizing only the most related primitives to the target concept.
To customize a human avatar, previous text-driven 3D avatar editing methods [24, 40, 9] are limited in terms of precision of concepts. To overcome this limitation, this paper aims to achieve controllable manipulation of 3D avatars towards a target concept in a convenient manner, akin to sliding a knob along a slider bar. However, achieving this goal is challenging due to the reasons outlined in the introduction. To address these challenges, we propose our Avatar Concept Slider (ACS), discussed in the following subsections. In Sec. 4.1, we address challenge 1 by leveraging LDA and devising a sliding loss to enable slider-like precise control of the editing process. To tackle challenge 2, discussed in Sec. 4.2, we augment the sliding loss with an attribute-preserving loss based on PCA, ensuring concept-specific editing that is disentangled from the identifying attributes to be preserved. Lastly, in Sec. 4.3, we introduce a concept-sensitivity-based primitive selection mechanism, facilitating faster editing via selective optimization of our avatar’s 3DGS primitives. Overall, our pipeline is built upon SDS optimization (introduced in Sec. 3), and our full method is illustrated in Fig. 2.
@@ 4.1 Precise Slider-like Concept Editing
To precisely control the 3D avatar editing towards a specific degree of the target concept, the first challenge is to identify the continuous and discriminative axes connecting two opposing extremes of the concept. This can be challenging because the concepts are often abstract and high-level, making it difficult to pinpoint where exactly these opposing sides and their connecting axes reside in the large feature space. To overcome this, we leverage LDA to identify the continuous concept axes between the user-provided pair of descriptions for the two opposing sides, by directly computing the axes that are the most discriminative to link the opposing sides within the large feature space. Then, based on the identified concept axes, we introduce a Concept Sliding Loss ($\mathcal{L}_{\text{CS}}$) to fine-tune an adapter, which enables controllable and precise concept-specific editing via SDS optimization (Sec. 3). Below, we first describe how we identify the concept axes, before discussing the Concept Sliding Loss.
@@
Concept Axes Identification with LDA. First, we aim to identify the continuous concept axes between the user-provided descriptions that represent two opposing sides of the target concept, where each point on the axis represents an intermediate degree of expression of the concept. By identifying the concept axes, we can precisely control the transitioning of the concept from one end to the other, and facilitate the learning of sliding capability via the sliding loss, enabling users to select any intermediate point along the concept axes for precise editing. Yet, this task is not straightforward. For instance, a simple solution may be to interpolate the semantic embeddings of two concept extremes, but such simple interpolation operations in a high-dimensional space often incurs high information redundancy and concept entanglement, incorporating many other unrelated concepts.
@@
Therefore, to handle this challenge, we leverage LDA. LDA [11, 23] is a statistical technique that aims to identify the most discriminative axes that well-separate different groups/classes of data samples, even while using a low dimensional representation. Thus, by employing LDA, we can directly compute a set of low-rank axes that are the most discriminative and crucial to link the opposing concepts (and their corresponding sets of features) within the large feature space. Specifically, we harness LDA to exploit the latent space of the pre-trained text-to-image diffusion model used during SDS optimization (Sec. 3) to facilitate understanding of the target concept in a high-quality latent space.
@@
Here, we first present the details of how we identify the concept axes. For simplicity, we explain it in the context of two-class LDA – where only positive and negative prompts are provided, although we can potentially use more classes (explained at the end of Sec. 4.1). In this case, at the beginning, the user provides a pair of text prompts representing the positive ($c_{p}$) and negative ($c_{n}$) sides of a concept. Then, based on these prompts, we sample two sets of latent features from the diffusion model $\boldsymbol{\epsilon}_{\boldsymbol{\phi}}$ by running the generation process $N_{s}$ times for each of the positive and negative sides, where we perform sampling over different diffusion timesteps using different seeds. This yields two sets of features $\{\mathbf{F}_{p,i,x}\}_{i\in[1,N_{s}],x\in[1,C]}$ and $\{\mathbf{F}_{n,i,x}\}_{i\in[1,N_{s}],x\in[1,C]}$ corresponding to the positive and negative sides, where $C$ refers to the channel dimensions of latent features. Each feature vector is $D$-dimensional, i.e., $\mathbf{F}_{p,i,x}\in\mathbb{R}^{D}$, $\mathbf{F}_{n,i,x}\in\mathbb{R}^{D}$.
@@
We then start the LDA analysis [11, 23] using these two feature sets for the two opposing sides of the target concept. Specifically, we first compute both the within-class scatter ($\mathbf{S}_{w}\in\mathbb{R}^{D\times D}$) and the between-class scatter ($\mathbf{S}_{b}\in\mathbb{R}^{D\times D}$), where $\mathbf{S}_{w}$ quantifies how much individual feature vectors deviate from their corresponding set centroids, while $\mathbf{S}_{b}$ measures the deviation between the centroids of both sides of the concept (more details in Supplementary). Then, using $\mathbf{S}_{b}$ and $\mathbf{S}_{w}$, the goal of LDA is to find concept axis $\mathbf{b}_{c}\in\mathbb{R}^{D}$, such that the projection of each feature vector onto the concept axis $\mathbf{b}_{c}$ leads to maximal ratio of $\mathbf{S}_{b}$ to $\mathbf{S}_{w}$. Intuitively, this encourages the feature vectors from both sets to be separated and positioned far from each other along the concept axis, while the feature vectors from the same set become tightly clustered along the concept axis. By doing so, LDA effectively pinpoints the direction of a concept axis, along which the opposing ends of the concept reside. Specifically, this process can be expressed as the following optimization problem:
@@
where the numerator ($\mathbf{w}\mathbf{S}_{b}\mathbf{w}^{T}$) and denominator ($\mathbf{w}\mathbf{S}_{w}\mathbf{w}^{T}$) on the right hand side of Eq. 2 represent the projected scatters over the concept axis, and we want to find a concept axis $\mathbf{b}_{c}$ that maximises their ratio while constraining $\mathbf{w}$ to be a unit vector. Notably, the solution to Eq. 2 can be conveniently computed as the leading eigenvector of $\mathbf{S}_{w}^{-1}\mathbf{S}_{b}$ (refer to Supp). We highlight that these concept axes $\mathbf{b}_{c}$ which are found by LDA are the most discriminative low-rank axes which best separates the positive and negative classes.
In practice, we can use more than 2 classes/groups of features. For instance, we can include three classes: positive, neutral, and negative, allowing us to compute two concept axes via LDA that hold strong discriminative power between these classes in the latent space, which are orthogonal to each other. We can then apply our sliding loss onto both concept axes for a stronger effect. Note that, even more classes can be included (with more prompts), but we found that using 3 classes is usually sufficient to obtain good results in practice.
@@
Concept Sliding Loss. The above process identifies the concept axes which represents the continuous transition between the two opposing sides of the target concept. Based on the identified concept axes, we introduce a Concept Sliding Loss to fine-tune an adapter to facilitate concept-specific editing of the 3D avatar. Since we adopt the SDS optimization pipeline (Sec. 3) to edit the 3D avatar via a pre-trained diffusion model, we aim to incorporate such concept-specific slider-like editing abilities into it. Hence, we propose a Concept Sliding Loss ($\mathcal{L}_{CS}$) to impart such concept transitioning abilities to the adapter, enabling us to precisely control the editing through a scale factor $\alpha$ in Eq. 1.
@@
Intuitively, our Concept Sliding Loss $\mathcal{L}_{CS}$ facilitates the controlled movements of features along the concept axis towards a specified intermediate point, which is specified by the user as a slider interpolation factor. Specifically, to control such movements of features during editing, our loss fine-tunes the adapter ($\Delta\phi$) such that the scale factor $\alpha$ in Eq. 1 gains the ability as the slider interpolation factor. Thus, at deployment, with the user-provided interpolation factor $\hat{\alpha}$, we can set it as the scale factor (i.e., setting $\alpha=\hat{\alpha}$), thus allowing the adapted diffusion model to precisely locate the desired intermediate point on the concept axes. The Concept Sliding Loss $\mathcal{L}_{CS}$, when using only positive and negative concept prompts, is defined as: